OpenStudy (anonymous):

Which expressions have a value of 18? Choose all answers that are correct. A. –3 – 21 B. 7 – (–25) C. 34 – 16 D. –4 – (–22) i got the first answer

OpenStudy (anonymous):

a i think is right but idk tho

OpenStudy (anonymous):

C

OpenStudy (anonymous):

and D

OpenStudy (igreen):

-3 - 21 = -24 7 + 25 = 32 34 - 16 = 18 -4 + 22 = 18

OpenStudy (igreen):

So your answers are C and D.

OpenStudy (anonymous):

–3 – 21= (-3) + (-21) So A is no-no

OpenStudy (anonymous):

thx guys